OpenStudy (anonymous):

(Calculus - Optimisation) Attempted this twice, but I couldn't get it :( A fuel tank is being designed to contain 200 cubic meters of gasoline. The design of the tank calls for a cylindrical part in the middle, with hemispheres at each end. The maximum length of a tank (measured from the tips of each hemisphere) that can be safely transported is 16 m. If the hemispheres are twice as expensive per unit of area as the cylindrical part, find the radius and height of the cylindrical part so that the cost of manufacturing will be minimal.

OpenStudy (anonymous):

V = pi*r^2*h = 200 cm^3 V = pi*r^2*16 = 200 cm^3 r = 1.99

OpenStudy (anonymous):

Now that is not the answer obviously.

OpenStudy (anonymous):

Minimize: SA = 2πr^2 + 2πrh h = 16

OpenStudy (anonymous):

volume of cylinder would be (16-2r)πr^2 16-2r is the h of that part you with me so far?

OpenStudy (anonymous):

Where did you get the 16-2r?

OpenStudy (anonymous):

or is that just the height for the cylinder, because i use r when i refer to the radius

OpenStudy (anonymous):

it is the height of the cylinder not including the hemispheres at the end

OpenStudy (anonymous):

Yes, okay I get it

OpenStudy (anonymous):

give me 5 minutes and i will be able to explain in full

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so this question is all about minimizing cost right? so we will need to create a function for cost, and then see what variable minimize that function

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

we are assuming that we are in a magic math world and the walls of the tank are infinitely thin, so we treat them by their area

OpenStudy (anonymous):

as the problem tells us to

OpenStudy (anonymous):

so first we need to be familiar with 2 equations of surface area. the first surface area of a cylinder is 2πrh. the surface are of a sphere is 4πr. we have to hemispheres so we can just combine them and think about them as a sphere. still with me?

OpenStudy (anonymous):

i mean 2 hemispheres (not to)

OpenStudy (anonymous):

Isn't the surface area of a cylinder 2πrh + 2πr^2 Hemisphere: 4πr^2

OpenStudy (anonymous):

you are right about the sphere (typo on my end). but the equation for the cylinder is assuming that the cylinder has ends. here the cylinder is open ended

OpenStudy (anonymous):

Oh, I see

OpenStudy (anonymous):

so here is the equation for cost. we will come back to it in a minute. remember that the hemispheres will be twice as expensive so their cost per area will be multiplied by 2 cost = 2(4πr^2) + 2πrh

OpenStudy (anonymous):

brb (2 minutes)

OpenStudy (anonymous):

Okay, got it

OpenStudy (anonymous):

forget what i said about h=16-2r. thats not necessarily true. h< or equal to 16-2r. now that we have our equation for cost, lets look at out equation for the h and r with respect to the volume

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

the volume is 200cm^3 so the equation we get is this: volume of hemishpheres+volume of cylinder = 200 4/3πr^3 + πr^2h = 200

OpenStudy (anonymous):

Right, got that

OpenStudy (anonymous):

go on ahead, i'll brb

OpenStudy (anonymous):

so now we can isolate r or h from the equation i just wrote and plug it in to the other equation for cost

OpenStudy (anonymous):

you're doing this in a calculus course right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So we isolate h from the equation for volume, put that into the cost equation and then find the derivative?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

keep in mind we may find multiple solutions where derivative =0. so first we need to make sure that the results of a given solution satisfy the parameters that h+2R< OR =16. another parameter would be the h and r are both positive

OpenStudy (anonymous):

so...have worked out the derivative yet?

OpenStudy (anonymous):

you are going to need the product and quotient rule. it might take a while

OpenStudy (anonymous):

Okay, i understand now

OpenStudy (anonymous):

I am just heading out to school actually, so I shall access the openstudy site from there

OpenStudy (anonymous):

You just saved my life, thank you so much :D

OpenStudy (anonymous):

did it work?

OpenStudy (anonymous):

didn't have time to calculate, I will contact you when I have though

OpenStudy (anonymous):

or you online later on?

OpenStudy (anonymous):

*are

OpenStudy (anonymous):

maybe, im actually in israel so im a few hours a head of where ever you are

OpenStudy (anonymous):

you in highschool?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

what time do u think youll need help?

OpenStudy (anonymous):

I'll be on in an hour/ an hour and 15 minutes

OpenStudy (anonymous):

i should be here. i need to go walk my dog

OpenStudy (anonymous):

Okay, thanks

OpenStudy (phi):